For a normal distribution, find the z-score that separates the highest 30% from the rest of the distribution. What formula do I use to get the answer and where to find the values to use in the formula?
Answer #1

solution ! - Given that BC 272) = 30% El-P (z z)=0.30 = P(Z <2) = 1-6.30 = P(zez) =0.7 = P(220. 5244) =0.7 z =0.5244